Calculate the expected rate of return for each of the following:       -      ...

  1. Calculate the expected rate of return for each of the following:
          -       A.       B.       C.       D.       E.   

    D1 = $0.90, g (which is constant) = 9.9%, and P0=$38.39

          -       A.       B.       C.       D.       E.   

    D1 = $0.30, g (which is constant) = 4.1%, and P0=$36.30

          -       A.       B.       C.       D.       E.   

    D1 = $0.65, g (which is constant) = 8.0%, and P0=$23.68

          -       A.       B.       C.       D.       E.   

    D1 = $1.80, g (which is constant) = 4.8%, and P0=$37.22

          -       A.       B.       C.       D.       E.   

    D1 = $1.40, g (which is constant) = 9.6%, and P0=$33.06

    A.

    9.6%

    B.

    13.8%

    C.

    12.2%

    D.

    4.9%

    E.

    10.7%

1 points   

QUESTION 9

  1. If D1 = $1.82, g (which is constant) = 10%, and P0 = $74.29, what is the stock’s expected dividend yield for the coming year? (Show your answer in decimal form to three places)

1 points   

QUESTION 10

  1. ABC Incorporated just paid a dividend of $1.92 on its common stock. Joe forecasts that the dividends for ABC will grow at a constant rate of 8% over the foreseeable future. Assuming that he is correct, how many total dollars of dividends should Joe expect to receive over the next five years? (Show your answer to the penny)
Answer #1

HI

As per policy we will solve only top most question here.

1. D1 = $0.90

g = 9.9%

P0 = $38.39

rate of return =k

as per Dividend discount model

P0 = D1/(k-g)

k-g = D1/P0

k = D1/P0 + g

k = 0.90/38.39 + 0.099

k = 0.122444

k = 12.24%

Hence option c is correct

same way

2. k = 0.3/36.30 + 0.041

k = 0.049264

k = 4.93%

hence option D is correct

3. k = 0.65/23.68 +0.08

k =10.75%

Hence option E is correct

4. k = 1.80/37.22 + 0.048

k =  9.6%

Hence option A is correct

5. k = 1.4/33.06 + 0.096

k = 13.8%

Hence option B is correct

Thanks